OpenStudy (anonymous):

What is the landing velocity of an object that is thrown vertically down with a velocity of 5m/s from a height of 25m?

OpenStudy (anonymous):

\[v^2=v_0^2+2gy \\v_0=5;y=25 \]

OpenStudy (anonymous):

22.69 m/s :D

OpenStudy (anonymous):

initial velocity= Vi= 5m/s height=h=25m final velocity=Vf=? using third equation of motion 2gh= Vf^2 -Vi^2 2(9.8)(25)= Vf^2 - 25 490= Vf^2 - 25 Vf^2= 515 Vf= sqrt (515) Vf= 22.69 m/sec